Two cells of EMF 2E and E and internal resistance 2r and r respectively are connected in parallel. Obtain the expression for the equivalent EMF and the internal resistance of the combination??

Grade:12

1 Answers

Susmita
5 years ago
Equivalent internal resistance is simply
req=r1r2/(r1+r2)=2r/3
For parallel connection total current
I=I1+I2
Or,Eeq/req=E1/r1+E2/r2
Or,Eeq=reqE1/r1+reqE2/r2=(r2E1+r1E2)/(r1+r2)=4E/3